OpenStudy (anonymous):

Identify the 13th term of the geometric sequence 10, 8, 6.4 …

OpenStudy (zepp):

The common ratio (r) for this is 8/10 = 6.4/8 = 0.8 First term is 10 (a) And the formula to find the xth term is \[\huge x_n=ar^{n-1}\]

OpenStudy (anonymous):

i dont know how to use the formula

OpenStudy (zepp):

\[\huge x_{15} = 10*0.8^{15-1} \approx 0.43\]

OpenStudy (anonymous):

a13 ≈ 0.01 a13 ≈ 0.02 a13 ≈ 0.55 a13 ≈ 0.68 but it has to be one of these :3

OpenStudy (zepp):

Ohh, it's 13th, I'm sorry, I found the 15th :P

OpenStudy (anonymous):

lol its okay :)

OpenStudy (zepp):

The one I gave you is the general formula to find pretty much any term in your sequence The # is the 'n'th term you are looking for so you plug 13 in (Not 15 haha). \[\huge x_{13} = ar^{13-1}\] Then a = 10, the first term r = 0.8, common ratio \[\huge x_{13} = 10 * 0.8^{13-1}\] Solve: \[\huge x_{13} = 10 * 0.8^{13-1} \approx 0.68\]

OpenStudy (anonymous):

thank you